OpenStudy (anonymous):

Multiply (1.2 ⋅ 1028) ⋅ (3 ⋅ 10−19). Express the answer in scientific notation. 3.6 ⋅ 109 3.6 ⋅ 1010 36 ⋅ 109 36 ⋅ 1010

OpenStudy (ryanwhite499):

(1.2)(1028)((3)(10)−19) =1233.6((3)(10)−19) =1233.6(30−19) =(1233.6)(11) =13569.6 Now look at your answers. Which one is similar to 13569.6 or close to

OpenStudy (ryanwhite499):

@pandaluvs

OpenStudy (anonymous):

im thinking, one sec.

OpenStudy (anonymous):

ummm i think it is D? im not sure....

OpenStudy (ryanwhite499):

Is there any exponents in this

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

one sec

OpenStudy (ryanwhite499):

Yes see. I could of gave you the wrong answer

OpenStudy (anonymous):

sorry.

OpenStudy (ryanwhite499):

3600000000 is the answer if this was solve

OpenStudy (ryanwhite499):

3.6(109) =3.6*109 =3.6*(10*10*10*10*10*10*10*10*10) =3600000000

OpenStudy (ryanwhite499):

So your answer would be A

OpenStudy (anonymous):

yes,

OpenStudy (ryanwhite499):

3.6 *10^9

OpenStudy (anonymous):

thanks so much!